OpenStudy (anonymous):

4+5e^(-x)=9 I think there is no solution though am I right?

OpenStudy (zarkon):

obvious solution ;)

OpenStudy (anonymous):

subtract 4 divide by 5 rewrite in equivalent log form change the sign

OpenStudy (anonymous):

or as zarkon said. make an educated guess

myininaya (myininaya):

\[4+5e^{-x}=9\] \[4-4+5e^{-x}=9-4\] \[5e^{-x}=5\] \[\frac{5}{5}e^{-x}=\frac{5}{5}\] \[1e^{-x}=1\] \[e^{-x}=1\] x=0

OpenStudy (anonymous):

\[1e^{-x}\] priceless

OpenStudy (zarkon):

lol

myininaya (myininaya):

i was trying to be like jim

OpenStudy (anonymous):

mavis beacon teaches typing

OpenStudy (anonymous):

4) 4+5e^-x=9 4 + 5e^-x = 0 5e^-x = -4 5 = -4e^x e^x = -5/4 write it as x = ln(-5/4) <--- when you put this into calculator, you will get an error Answer: No solution

OpenStudy (anonymous):

where did the nine go?

myininaya (myininaya):

what ! 4+5e^-x=9 4 + 5e^-x = 0 <--- what happen to 9? 5e^-x = -4 5 = -4e^x e^x = -5/4

OpenStudy (anonymous):

good ?

OpenStudy (anonymous):

nine of one, three quarters of a dozen of the other

myininaya (myininaya):

no not good

OpenStudy (anonymous):

This is what I got when I punched it into web 2.0

myininaya (myininaya):

what happen to 9?

myininaya (myininaya):

how does 9=0?

OpenStudy (anonymous):

This is why I asked the question?

myininaya (myininaya):

oh so this isn't your work?

myininaya (myininaya):

i don't know what web 2.0 is

OpenStudy (anonymous):

No it is web 2.0 work not mine.

myininaya (myininaya):

don't use it its horrible if it did that

OpenStudy (anonymous):

It is a scientific calculator on the web.

OpenStudy (anonymous):

use wolfram alpha or geogebra

myininaya (myininaya):

try this: wolfram

myininaya (myininaya):

http://www.wolframalpha.com/

myininaya (myininaya):

i think it also shows you the work sometimes